OpenStudy (anonymous):

Find 1st term and common difference: 2,7,12,17

OpenStudy (saifoo.khan):

common difference = 5

OpenStudy (saifoo.khan):

1st term is 2

OpenStudy (anonymous):

7-2=12-7=17-12=5